A monopolist faces the demand curve Q = 120 - P/3. The cost function is C = Q 3 . Find the output that maximises this monopolist’s profits. What are the prices at profits and that output? Find the elasticity of demand at the profit maximising output.


1
Expert's answer
2021-03-08T09:22:20-0500

Let's first write the inverse demand function:


"P=360-3Q."

By the definition of the total revenue, we have:


"TR=PQ=(360-3Q)Q=360Q-3Q^2."

Then, we can find the marginal revenue:


"MR=\dfrac{dTR}{dQ}=360-6Q."

Let's find the marginal cost:


"MC=\dfrac{dTC}{dQ}=\dfrac{d}{dQ}(Q^3)=3Q^2."

The monopolist maximises its profit when MC=MR:


"3Q^2=360-6Q,""Q^2+2Q-120=0."

This quadratic equation has two roots: "Q_1=10" and "Q_2=-12". Since the quantity produced can't be negative the correct answer is "Q=10".

Then, we can find the price that maximises the profit:


"P=360-3Q=360-3\cdot10=\$330."

We can find the price elasticity of demand as follows:


"\epsilon=-\dfrac{P}{Q}\dfrac{dQ}{dP},""Q=\dfrac{P-360}{3}, \dfrac{dQ}{dP}=-\dfrac{1}{3},""\epsilon=-\dfrac{3P}{P-360}\cdot-(\dfrac{1}{3})=\dfrac{P}{P-360}."

The price elasticity of demand when P = $330 equals:


"\epsilon=\dfrac{\$330}{\$330-360}=-11."

Therefore, demand is elastic.